how do i work out x+11=y and 7x-49=y? (two differwnt equations)
Answer by nyc_function(2741)

Solve for x in x + 11 = y.
x + 11 = y
x = y - 11
Now plug that into the other equation to find the value of y.
7x - 49 = y
7(y - 11) - 49 = y
7y - 77 - 49 = y
-126 = -7y + y
-126 = -6y
-126/-6 = y
21 = y
We lastly replace y with 21 in EITHER of the equations given to you to find x.
I will use x + 11 = y but you could use 7x-49 = y if you want to.
x + 11 = y
x + 11 = 21
x = 21 - 11
x = 10
Done!